OpenStudy (anonymous):

Simplify the expression 8 − 3n when n = 12. Select one of the options below as your answer: * A. -28 * B. 28 * C. 18 * D. -18

OpenStudy (smurfy14):

A

OpenStudy (anonymous):

thank you (: ill reward you <3

OpenStudy (anonymous):

A -28

OpenStudy (anonymous):

Plug n=12 in \(8-3n\), you get \(8-3(12)=8-36=-28\).

OpenStudy (smurfy14):

youre welcome and thanks :)

OpenStudy (anonymous):

thank u so much i hate plato